by how much would the 90 billion 2008 tax rebates have shifted AD if the MPC was 0.95?
Answer #1


Tax rebate leads to increase in disposable income.

Increase in disposable income leads to increase in consumption which in result leads to increase in aggregate demand (AD).

Tax rebate = $90 billion

MPC = 0.95

Disposable income will increase by $90 billion

Increase in consumption = Disposable income * MPC

Increase in consumption = $90 billion * 0.95 = $85.5 billion

Calculate Multiplier -

Multiplier = 1/(1-MPC)

Multiplier = 1/(1-0.95)

Multiplier = 1/0.05

Multiplier = 20

Calculate the increase in AD -

Increase in AD = Increase in consumption * Multiplier = $85.5 billion * 20 = $1,710 billion

Thus,

The AD will shift by $1,710 billion.